Commissioning Notes from 9th-10th June 2008 MST

Observers: A.Rakich, D.Miller, J.M. Hill, T. Leibold, N. Cushing and S. Hooper in Tucson. OSA: David G.-H.

Summary

Lost time to a few problems early on as discussed below. Once we got on sky we found a problem with the offset-to-hotspot command going the wrong way, the offset seemed to be of the right magnitude and wrong direction. We overcame this by manually pointing the star to within half a guiding box sized box, so that it was still inside the box after the intial offset, then guiding took over and guiding worked well. We had not a single unexpected GCS failure, oacserver error, RPC communication error or anything else of that ilk through the night, and stayed guiding on the same star for several hours taking data to chase down the 500 nm focus problem. At the end of the night we went back to an off-axis star and confirmed that the star was where it should be, our only problem with getting guiding going at the moment is the offset-to-hotspot having the wrong sign for correction.

Night time Activities

(Times in MST below)

Sundown - 12:00: chased a few problems, the most annoying one being that the M2 cover was left on and we wasted a couple of hours before we worked that out. Finally point to a star and it is within ~15" of where it should have been and not extremely aberrated.

Start attempting to guide with new build of GCS.

12:05 Achieved guiding on-axis. Torseten is happy as this verifies various signs went in the right way. Going to off-axis star.

12:32 PCS and Dave T's "point-to" script seemingly disagree with the coordinates of an off-axis star. For a star 218" off-axis, PCS was giving XY ccordinates for the AGW stage of (0, 480), and Dave's Y number was 603. BS9177 #19 at default PA of -43.

1:42 /Repository/AGW_data/080610w for WFS images. First one in sequence is WFSC 00006.

12:42 On-axis guiding now not working either.

1:00-1:10 BS9177 on-axis. Many failed offset-to-hotspot attempts here. Trying to send star to hotspot. Seems that the offset to the hotspot is of the wrong sign.

1:14 Norm and Steve leave, promising that software will work hard on all these problems tomorrow. (today). We perservere to see if we can revive on-axis guiding

1:25 Achieved on-axis guiding by manually pointing the star right onto the hotspot. Once star is in guide box guiding proceeds well.

1:35 Starting up wavefront sensing using Doug's IDL routine. First order of business will be to do more measurements to track down the "500 nm extra focus" issue we found in May. For this we wil repeat our tests driving in 1000 nm of Z4, Z5, Z7 and Z11. This time we will then reverse sign and make new sets of measurements without first running active optics.

1:48 WFSC00007 first set of corrections sent to primary.

1:53 changed the sign of the Z4 coefficient in Doug's IDL routine to take into account Chris's recent change of Z4 sign. Image quality is ~0.5 arcsec on the guider images (1 sec exposure).

2:01 rotator got to dead zone (350-360 degrees) about ten minutes ago. Now have to reset PA to say 180 degrees, and manually point the star onto the hotspot (294, 252).

2:16 WFSC 000016-20 are our first five WFS measurements set at zero aberration.

2:20 issue with guiding. We see an image every ~2 seconds from the guide camera, even though the guide rate is set to 0.4 seconds.

2:21 set Z4 to +1000 in PSF. Take new set of five WFSC shots WFSC000021-25

2:25 set Z4 -1000 nm (back to zero). WFSC 000026-30

2:30 set Z4 -1000 WFSC 000031-35

2:33 send Z4 +1000 re-collimate

2:38 WFSC 000039-43 first set of five images for Z5 block

2:41 clear coefficients in PSF, send +1000 Z5, WFSC 000044-48

2:46 clear coefficients in PSF, send -1000 Z5 (back to zero) WFSC000049-53

2:50 send -1000 Z5 WFSC 54-58

2:54 back to zero and re-collimate

3:01 Had to roll the PM off an x-limit using global offsets. 0.5 mm -X and +10" RY. Were able to keep guiding lock by doing 0.5 mm DX and 5" RY first (pointing free adjustment), then stepping in 1" steps for the other 5" to give the coma free adjustments.

3:02 recollimate with ActO

3:05 first set of recollimated data at WFSC000064-68 (for the delta Z7 run).

3:09 Z7 +1000 WFSC000069-73

3:13 Z7-1000 (back to zero) WFSC000074-78

3:17 Z7-1000 WFSC 79-83

3:22 back to zero and collimate

3:25 collimated and doing run of 5 @ zero for Z11 run WFSC 87-91

3:29 clear coefficients add +1000 Z11 WFSC000092-96 focus jumps a large amount (from around zero to around +1600 nm).

3:34 -1000 Z11 WFSC000097-101

3:40: -1000Z11 WFSC000102-106

3:48 recollimate

3:50 will look at this lost motion in focus business. After collimating take a sequence of sets of five measurements with focus set to 0nm, 200nm, 400nm, 600nm first set of 5 measuremnts at zero, WFSC000109-113

3:55 WFSC000114-118 +200 Z4

3:59 WFSC000119-123 +200 Z4 (total +400)

4:03 WFSC000124-128 +200 Z4 (total +600) (seem to have a very much slower PSF GUI display, has the 4:00 am problem come back?)7

4:07 WFSC000129-133 +200 Z4 (total +800) (GUI more responsive now)

4:11 WFSC000134-138 -200 Z4 (total +600)

4:15 WFSC000139-143 (total +400)

4:19 recollimate preparing for run of Z5 incremental measurements

4:20 hit the +X limit again. repeated excercise described above to roll in anbother 0.5 mm of -Dx

4:23 sunlight starting to affect WFSC images, will go back to star # 19 and see if we can see it on the full chip view of the guide camera

4:28 found star on guide camera, getting washed out now, calling it a night.

-- AndrewRakich - 10 Jun 2008
Topic revision: r2 - 11 Jun 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback